### v1

# Two Heads Are Better Than One: A Multi-Agent System Has the Potential to Improve Scientific Idea Generation

Abstract
The rapid advancement of scientific progress requires innovative tools that can accelerate discovery. While recent AI methods, particularly large language models (LLMs), have shown promise in tasks such as hypothesis generation and experimental design, they fall short in replicating the collaborative nature of real-world scientific practices, where diverse teams of experts work together to tackle complex problems. To address the limitation, we propose an LLM-based multi-agent system, i.e., Virtual Scientists (VirSci), designed to mimic the teamwork inherent in scientific research. VirSci organizes a team of agents to collaboratively generate, evaluate, and refine research ideas. Through comprehensive experiments, we demonstrate that this multi-agent approach outperforms the state-of-the-art method in producing novel and impactful scientific ideas, showing potential in aligning with key insights in the Science of Science field. Our findings suggest that integrating collaborative agents can lead to more innovative scientific outputs, offering a robust system for autonomous scientific discovery.
1 Introduction
The rapid acceleration of scientific progress necessitates more efficient and innovative tools for exploring new concepts and tackling complex challenges (Park et al., 2023b). The concept of automatic scientific discovery has emerged as a promising solution to expedite innovation, representing a long-standing ultimate goal within the scientific community (Langley, 1987). With the development of artificial intelligence (AI), automatic scientific discovery has witnessed the potential to revolutionize how research is conducted by automating key steps in the scientific process, ranging from hypothesis generation to experimental design (Raghu & Schmidt, 2020; Spangler et al., 2014).
Figure 1: The proposed LLM-based multi-agent system, VirSci, consists of five key steps: Collaborator Selection, where a research team is assembled; Topic Discussion, where the research topic is determined; Idea Generation, where team members propose and refine ideas; Novelty Assessment, where ideas are evaluated and voted on to select the best one; and Abstract Generation, where the selected idea is developed into a complete abstract.
More recently, foundational models, especially the large language models (LLMs) (OpenAI, 2023; Dubey et al., 2024), have shown significant progress in general capabilities, facilitating their applications in various stages of scientific discovery (Abramson et al., 2024; Chang & Ye, 2024), including literature reviews (Hsu et al., 2024), experimental designs (Huang et al., 2024), etc. A notable development is the AI Scientist (Lu et al., 2024), which introduces a scalable system for end-to-end scientific paper generation, highlighting the potential of LLMs to drive autonomous scientific discovery. Despite its capabilities, AI Scientist operates with ONE agent, which falls short of replicating real-world scientific practices, where research is often conducted through collaborative efforts involving diverse teams of experts (Kayacik et al., 2019).
To address the limitations of a single executive system, such as the AI Scientist, and to better replicate the collaborative nature of real-world scientific discovery (Gauch, 2003), we focus on the idea generation phase in the research process, which demonstrates more collaborative aspects (Linsey et al., 2005). From this end, we propose an LLM-based multi-agent system, Virtual Scientists (VirSci), designed to harness the potential of LLM agents in assisting autonomous scientific idea generation. Leveraging the inherent human-like reasoning capabilities of LLMs (Xie et al., 2024), VirSci simulates the collaborative process of research idea generation, which is divided into five steps (see Fig. 1): (1) Collaborator Selection, (2) Topic Selection, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To be more specific, we construct a knowledge bank of the background of scientists of interest and develop digital twin (Tao et al., 2018) agents of real scientists using a retrieval-augmented generation (RAG) framework (Gao et al., 2023). The lead agent, or team leader, identifies appropriate collaborators based on the scientist collaboration network, aligning their expertise and research interests to reflect real-world cooperation patterns. Using a past paper database, the team retrieves reference works to guide the formation of novel ideas. Moreover, we implement a “team discussion” mechanism in subsequent stages, where collaborators engage in iterative inter- and intra-refinement dialogues, to enhance the quality of each output. Finally, the team generates a comprehensive abstract representing the proposed ideas. To evaluate the novelty of the ideas, we introduce a benchmark for the measurement from three perspectives: dissimilarity to past papers, alignment with contemporary research trends, and the potential influence in contemporary research (Shao et al., 2020; Yang et al., 2022). By comparing the abstract against both a past and a contemporary paper database, we ensure the generated ideas are innovative while aligning with emerging scientific directions, thus validating the effectiveness of our approach.
We conduct comprehensive experiments to verify the effectiveness of VirSci in producing novel scientific ideas. The findings demonstrate that the multi-agent system outperforms the single-agent executive with an average gain of 13.8% and 44.1% in alignment level and potential impact on contemporary research, respectively. Additionally, our experiments reveal emergent social behaviors among agents, aligning with prior studies in Science of Science (Fortunato et al., 2018; Wu et al., 2019; Zeng et al., 2021; Shi & Evans, 2023), suggesting the potential for further exploration of the mechanisms in research collaboration using multi-agent simulations. To sum up, our core contributions are summarized as follows:
  • To the best of our knowledge, we propose the first multi-agent system for conducting scientific collaborations in an end-to-end pipeline from team organization to novel scientific idea generation. Furthermore, the real data is utilized for role-play and the objective evaluation of final outputs.
  • We conduct extensive evaluations to investigate VirSci in terms of the team settings and the novelty of generated scientific ideas. The results demonstrate that multi-agent collaboration can improve the quality of the outcomes, surpassing the SOTA single-agent method.
  • The simulation results align with the important findings in Science of Science, such as fresh teams tend to create more innovative research, showcasing the potential of VirSci as a powerful tool for future research in this field.
2 Related Work
2.1 AI for Scientific Discovery
In recent years, AI has fundamentally reshaped the landscape of scientific discovery by providing powerful tools that enhance various research processes (Xu et al., 2021). AI techniques, especially generative AI, can facilitate basic scientific discoveries, such as identifying complex molecular (Vignac et al., 2022) and protein structures (Abramson et al., 2024), drastically reducing the time required for experimental iterations. These advancements have found wide application across diverse fields such as chemistry (Liu et al., 2023a), meteorology (Bi et al., 2023), and medicine (Rajpurkar et al., 2022), etc. Besides, with the advent of LLMs, AI methodologies can step further and collaborate in streamlining critical stages of the scientific pipeline, including hypothesis generation, experimental design, data acquisition, and analysis(Zheng et al., 2023; Wang et al., 2023; Miret & Krishnan, 2024; Wysocki et al., 2024; Lu et al., 2024). Nevertheless, these approaches lack the collaborative nature of the scientists intrinsic to real-world research. VirSci is the first to harness the power of an LLM-based multi-agent system to facilitate the generation of research ideas in autonomous scientific discovery.
2.2 Multi-agent Systems in Team Collaboration
A multi-agent system for team collaboration leverages autonomous agents to coordinate, communicate, and solve tasks within a shared environment, mimicking human teamwork dynamics (Dorri et al., 2018). Traditional multi-agent systems typically involve semi-autonomous agents coordinating through explicit protocols and structured messages to achieve common goals (Dunin-Keplicz & Verbrugge, 2011; Bakliwal et al., 2018). The advent of LLMs has revolutionized this landscape by enabling agents to utilize natural language for communication and collaboration in a believable proxy of human behavior (Park et al., 2023a), thereby fostering a more intuitive and flexible interaction model. Recent studies have further verified the superior performance of LLM-based multi-agent systems in various domains, such as programming, game playing, and complex reasoning tasks when compared to single-agent execution (Liu et al., 2023b; Wang et al., 2024; Light et al., 2023; Du et al., 2024). In this work, we strategically implement the power of LLM-based multi-agent systems to function as collaborative scientists, promoting de novo scientific ideas.
3 The Virtual Scientists
Figure 2: Key components of the proposed system. The left section illustrates the collaborator selection process, where the team leader forms a research team. The middle section highlights the discussion routine, a fundamental part of every step in the system, where the team engages in collaborative dialogue to progress through tasks. The right section depicts the architecture of the author knowledge bank and paper database, which provide critical information used throughout the collaboration process.
In this paper, we aim to build a multi-agent system using a real-world academic dataset to simulate how a scientist assembles a research team and collaboratively generates an abstract that details a novel scientific idea. Our VirSci system consists of two components: a scientific research ecosystem and a multi-agent system for scientific idea generation.
3.1 The Scientific Research Ecosystem
The scientific research ecosystem comprises two main components: paper information and corresponding author information ranging from year $y_{\text{start}}$ to $y_{\text{end}}$. First, we select a year $y_{\text{bound}}$ as a time point and split the papers into two subsets: past papers $B_{\text{past}}$ and contemporary papers $B_{\text{con}}$. We further extract authors from $B_{\text{past}}$ to form the complete set of scientists $S$, with each scientist’s background information stored in the author knowledge bank, and the adjacency matrix $A$, which represents the collaboration counts between scientists.
Past Paper Database To construct the past paper database $B_{\text{past}}$ using the Faiss, we selected papers published before the $y_{bound}$. Each paper includes essential information such as its title, citation count, and abstract.
Contemporary Paper Database The contemporary paper database $B_{\text{con}}$, also constructed with Faiss, consists of papers published after $y_{bound}$. Similarly, each paper’s basic information is structured in the same way as the past papers. Although using papers from this time range may raise concerns about data leakage, given that LLMs are trained on data within this period, we will explain in detail why this does not pose a threat to the overall validity of our experiments and conclusions in the Appx. D.
Author Knowledge Bank  For each scientist in $S$, we extract their basic profile from the computer science dataset, which includes their name, affiliations, citation count, research interests, and collaboration history. Using the KnowledgeBank module from AgentScope(Gao et al., 2024), we embed these scientist profiles into the author knowledge bank. This allows agents to quickly access and familiarize themselves with other initialized agents’ information. Notably, real author names are masked to prevent data leakage and privacy problems during agent initialization (See Appx. C).
Adjacency Matrix Given the scientist set $S$, let $A$ represent the adjacency matrix, where $A_{i,j}$ denotes the number of times scientist $i$ has collaborated with scientist $j$. To prevent agents from always choosing previously collaborated scientists, overlooking the benefit of fresh collaborations that often lead to more original and impactful research (Zeng et al., 2021), we increment all values in $A$ by 1. This adjustment ensures that scientists with no prior collaborations still have a chance of being selected, encouraging agents to explore new partnerships.
3.2 The Multi-agent System for Scientific Collaboration
We first randomly sample a scientist $s_{0}$ from $S$ as the team leader. The team leader then follows these steps to produce an abstract: (1) Collaborator Selection, (2) Topic Discussion, (3) Idea Generation, (4) Idea Novelty Assessment, and (5) Abstract Generation. To help each agent become familiar with the backgrounds of other team members without overloading the initialization prompt, we employ retrieval-augmented generation (RAG) (Lewis et al., 2020), used throughout all five steps. All necessary prompts and example scenarios are shown in Appx. G and H.
Collaborator Selection The first step in our system is collaborator selection, aimed at forming a team of scientists, $T=\{s_{0},\dots,s_{i},\dots,s_{n}\}$, where $n$ denotes the team size. When $s_{0}$ is selecting collaborators, we convert the adjacency matrix, $A$, into a probability distribution using the following equation: $P_{i,j}=\frac{A_{i,j}}{\sum_{j=1}^{N}A_{i,j}}$, where $N$ denotes the size of $S$. This allows the team leader to iteratively send invitations to preferred collaborators. Upon receiving an invitation, the invited scientist evaluates whether to join the team using the chain-of-thought process (Wei et al., 2022), considering the profiles of $s_{0}$ and the current team members. If accepted, the scientist is added to the team $T$. This process continues until the pre-defined team size $n$ is reached.
Topic Discussion The next step is to propose a research topic, which will guide the research direction. Inspired by multi-round collaboration (Mezirow, 2003; Sunstein, 2005; Amgoud & Prade, 2009) and multi-agent collaboration strategies (Xu et al., 2023; Zhang et al., 2023; Shinn et al., 2024), we design a general team discussion mechanism. In this mechanism, team members engage in discussions based on a specific task description prompt. This process is applied not only to topic discussion but also to subsequent collaboration steps. While allowing agents to decide when to stop the discussion would better reflect real-world scenarios, fixing the number of turns ensures consistent inference costs across different team settings in our experiments. Therefore, we leave the discussion of adaptive turn numbers to the ablation study (See Sec. 5). Given the team $T$, the prompt during the topic discussion is:
$Q_{k,i}=\langle Q_{team},Q_{topic},\overset{k-1}{\underset{t=1}{\bigcup}}(\overline{D_{t}}),\overset{i-1}{\underset{j=0}{\bigcup}}(R_{k,j})\rangle,$ (1)
[…]
Idea Generation Third, the team is tasked with proposing several potential ideas. To align with genuine research workflows and mitigate LLM illusions (Huang et al., 2023), each agent is required to generate a comprehensive response that includes three key components: (1) a description of the idea, (2) a specific experimental plan, and (3) a self-assessment covering metrics such as novelty, feasibility, and clarity, representing the agent’s confidence (See Appnx. 12).
At the start of the idea generation process, when no ideas have yet been proposed, the agent is provided with references by searching $B_{past}$ using the topic $R_{topic}$, denoted as $B_{past}(R_{topic})$. The first idea-generation prompt is defined as:
$Q_{1,0}=\langle Q_{idea},R_{topic},B_{past}(R_{topic})\rangle,$ (2)
where $Q_{idea}$ represents the task description. Inspired by the concept of gradually expanding an archive of ideas (Zhang et al., 2023; Lu et al., 2024), when a scientist $s_{i}$ at turn $k$ receives an existing idea from the previous response $R_{k,i-1}$, we retain the previously generated ideas along with their corresponding references from $B_{past}$. These are passed to the next agent, who can either refine the existing idea or propose a new one, depending on its choice. The prompt is represented as:
$Q_{k,i}=\langle Q_{idea},R_{topic},B_{past}(R_{k,i-1}),\overset{k-1}{\underset{t=1}{\bigcup}}(\overline{D_{t}}),\overset{i-1}{\underset{j=0}{\bigcup}}(R_{k,j})\rangle.$ (3)
Afterwards, the response of $S_{i}$ at turn $k$ can be represented as $R_{k,i}\sim P_{s_{i}}(\cdot|Q_{k,i})$. After $K$ turns of discussion, we retain the three ideas with the highest confidence and store them in the idea list $I$.
Novelty Assessment To enhance the quality of ideas and mitigate agent overconfidence, we introduce an idea novelty assessment, enabling agents to compare each idea with related papers from $B_{past}$ and vote for the idea they consider most novel. Given the idea list $I$, agents search for related papers using each idea’s description to determine whether it significantly overlaps with existing works. To simulate a blind review process, no dialogue memory is included in the prompt. The prompt for $s_{i}$ at turn $k$ is defined as:
[…]
where $I_{j}$ is the $j$-th idea in $I$. Following the chain-of-thought process, the response $R_{k,i}\sim P_{s_{i}}(\cdot|Q_{k,i})$ includes the scientist’s preferred idea and the reasoning behind their choice. The idea receiving the highest number of votes is then selected as the final idea, $R_{idea}$, for abstract generation.
Abstract Generation Lastly, the team is required to produce a comprehensive abstract that includes the following sections: (1) Introduction, (2) Objective, (3) Methods, (4) Expected Results, and (5) Conclusion (Alexandrov & Hennerici, 2007). At the start of abstract generation, the team leader provides an initial draft based on $R_{idea}$. The first abstract-generation prompt is:
$Q_{1,0}=\langle Q_{abstract},R_{idea}\rangle,$ (5)
where $Q_{abstract}$ represents the task description and format requirements.
When an abstract is provided by the previous response $R_{k,i-1}$, the next scientist’s response should include: (1) an evaluation of the prior abstract (evaluation metrics are detailed in Appx. 15), (2) proposed modifications, and (3) the revised abstract to enable continuous refinement. The corresponding prompt is:
$Q_{k,i}=\langle Q_{abstract},Q_{judgement},R_{k,i-1}\rangle,$ (6)
where $Q_{judgement}$ is the prompt that asks agents to evaluate the previous abstract. Dialogue history is not included in this prompt since the process is iterative and focuses on refining a single abstract. Including previous versions would make the prompt redundant. After $K$ turns of revision, the final abstract is denoted as $R_{abstract}$.
A self-review mechanism is also considered after $R_{abstract}$ is finalized to pre-check its novelty. The optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$, where more details are shown in Appx. E.1. Because it introduces uncertainty in total inference cost, making it difficult to ensure fair experimental comparisons, we only discuss the effectiveness of this module in the ablation study (see Sec. 4.4).
4 Empirical Study
4.1 Experimental Settings
Dataset We build our scientific research ecosystem using real scientists’ information from the AMiner Computer Science Dataset  [^1] https://www.aminer.cn/aminernetwork, which was constructed by extracting scientists’ profiles from online web databases (Tang et al., 2008). This dataset includes 1,712,433 authors and 2,092,356 papers, covering the period from 1948 to 2014, with disambiguated author names. To manage the large volume of data, we set $y_{start}$, $y_{bound}$, and $y_{end}$ to 2000, 2010, and 2014, respectively. For quality assurance, we filtered out past papers lacking abstracts or with fewer than 10 citations, contemporary papers with fewer than 5 citations or missing abstracts, and authors with fewer than 50 papers or 50 co-authors. As a result, we extracted detailed information from 156 authors and 85,217 papers to construct the ecosystem and initialize the corresponding agents for the simulation. All paper and author data are embedded using the “mxbai-embed-large” model (Lee et al., 2024).
Implementation We implement our system on top of the Agentscope framework (Gao et al., 2024), which serves for LLM-empowered multi-agent applications. We evaluate our system using different publicly available LLMs: GPT-4o (OpenAI, 2023) and Llama-3.1 (8b and 70b) (Dubey et al., 2024). GPT-4o is accessible exclusively via a public API, while the Llama-3.1 models are open-weight and invoked using the Ollama (Ollama, 2024) in our experiments. Each experimental run on Llama-3.1 (8b) takes approximately 10 minutes on 1 NVIDIA A100 40G GPU within a team discussion setting of 4 members and 5 turns ($K$ = 5). All experimental results are averaged on 20 runs.
Evaluation Metrics Since no single evaluation metric perfectly captures the novelty of scientific outputs, we employ three common metrics that align with our intuition: (1) Historical Dissimilarity (HD): The average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in $B_{\text{past}}$ (Shao et al., 2020; Zhou et al., 2024). A larger distance indicates that the generated abstract is more dissimilar from existing papers, suggesting a higher likelihood of novelty. (2) Contemporary Dissimilarity (CD): The average Euclidean distance between the generated abstract embedding and embeddings of the top 5 most similar abstracts in $B_{\text{con}}$. A smaller distance indicates greater similarity to newer papers, also suggesting a higher likelihood of novelty. (3) Contemporary Impact (CI): The average citation count of the top 5 most similar abstracts in $B_{\text{con}}$ (Yang et al., 2022). A higher citation count suggests that the generated abstract is more likely to have a higher impact. To ensure comparability, we normalize each calculated metric using the mean value derived from the entire corresponding database, with normalization defined as the metric divided by the mean value. Since novelty is difficult to measure directly, we introduce a proxy metric to comprehensively account for the three indicators: (4) Overall Novelty (ON). We assume that ON is positively related to both HD and CI and negatively related to CD, calculated as $\text{ON}=(\text{HD}\times\text{CI})/{\text{CD}}$. Mathematically, the expected value of ON is proportional to the true novelty. Further discussion of this metric is provided in the Appx. F.
Experimental Design We aim to explore several key aspects of VirSci’s performance. (1) How do the evaluation metrics of abstracts from VirSci outperform those generated by the SOTA single-agent method AI Scientist under similar conditions (Sec. 4.2)? (2) How do the novelty scores of abstracts improve in relation to different team settings, including team size and discussion turns (Sec. 4.3.1), team freshness (Sec. 4.3.2), and team research diversity (Sec. 4.3.3)? (3) How do different components in our system affect the novelty of the generated abstracts (Sec. 4.4)? (4) How do adaptive turn numbers affect the novelty of the generated abstracts (Sec.5)?
4.2 Comparisons with AI scientist
Table 1: Comparisons with AI Scientist. Results show that our multi-agent system outperforms the AI Scientist across all metrics, with GPT-4o achieving the highest performance.
Agent Model Method LLM Review Score $\uparrow$ CD $\downarrow$ CI $\uparrow$
GPT-4o1 AI Scientist 3.10 0.38 3.22
Ours 3.34 (+0.24) 0.34 (-0.04) 3.78 (+0.58)
LLaMA3.1-8b AI Scientist 2.09 0.49 2.12
Ours 2.31 (+0.22) 0.42 (-0.07) 3.29 (+1.17)
LLaMA3.1-70b AI Scientist 2.24 0.48 2.11
Ours 2.53 (+0.29) 0.40 (-0.08) 3.36 (+1.25)
  • 1
    GPT-4o API is “gpt-4o-2024-08-06”.
Adjusting Settings for Fair Comparison with AI Scientist Given that the problem formulation of the AI Scientist differs from ours, we must make several justifications to ensure relative fairness in our comparisons: (1) Since the AI Scientist is limited to generating ideas from its pre-defined topics (2D Diffusion, NanoGPT, and Grokking), we include NanoGPT in the topic selection prompt for VirSci as the initial discussion topic, ensuring that the final abstracts align with the same research direction. (2) Since our teamwork approach differs from the AI Scientist’s solitary investigation, we need to ensure that the comparisons are made under the same inference cost. The AI Scientist performs 50 turns of self-reflection during its idea generation, which does not apply to its paper generation. To align the inference costs, we set the number of team members to 4 and the number of discussion turns to 5, ensuring the experiments are conducted under approximately the same computational cost. (3) Since the AI Scientist lacks a scientific research ecosystem, it retrieves papers across all time ranges through the Semantic Scholar API (Fricke, 2018). To maintain consistency with the AI Scientist, we replace our databases with the Semantic Scholar API for both paper retrieval and metric calculation. Specifically, after generating ideas and corresponding abstracts from both the AI Scientist and our system, we use the generated ideas as queries to retrieve related papers, extracting corresponding abstracts and citation counts for evaluation. (4) We evaluate the generated abstracts using both our metrics (CD and CI) and the AI Scientist’s metric (LLM review score) (Lu et al., 2024). The LLM review score is calculated by GPT-4o which conducts abstract reviews based on a truncated version of the Neural Information Processing Systems (NeurIPS) conference review guidelines [^2] https://neurips.cc/Conferences/2024/ReviewerGuidelines, shown in Fig. 17.
Experimental Results and Analysis As shown in Fig. 1, our multi-agent system outperforms the AI Scientist across all three metrics: CD, CI, and LLM review score. This demonstrates that our approach effectively enhances the novelty of generated abstracts in a collaborative setting. Notably, GPT-4o consistently achieves the highest novelty scores among the models tested, reflecting its superior ability to generate innovative ideas and abstracts. In contrast, the Llama-3.1 (8b and 70b) models do not show a significant difference in novelty scores, suggesting that moderate changes in model size may not enhance novelty.
4.3 Exploring Science of Science: The Impact of Team Dynamics on Novelty
While the effects of team size, team freshness, and team research diversity on the novelty of research outputs have been established in the Science of Science field using traditional statistical methods (Wu et al., 2019; Zeng et al., 2021; Shi & Evans, 2023), they have yet to be verified in an LLM-based multi-agent system. We conduct the following experiments to demonstrate VirSci’s potential to simulate key findings in Science of Science.
4.3.1 How Team Size and Discussion Turns Affect Novelty
Figure 3: Effects of team size and discussion turns on novelty scores. Peak innovation occurs with 8 members and 5 turns, while larger teams or excessive turns hinder creativity. “Inference Cost” is the product of team size and turns.
Effects of Team Size on Novelty The results shown in Fig. 3 indicate that increasing the number of team members can enhance the novelty score of the generated abstracts. By adding new team members, a broader range of ideas and perspectives can be facilitated, leading to more creative solutions and innovative outputs. However, this relationship is not strictly linear; our findings suggest that the peak novelty occurs with a team size of 8 members. While a moderate increase in team size may boost novelty, excessively large teams can introduce coordination challenges and communication barriers. These issues may dilute individual contributions and foster groupthink, where the focus shifts from original ideas to achieving consensus. Thus, there appears to be an optimal team size that maximizes creativity without overwhelming the collaborative process. This conclusion aligns with existing literature, which suggests that while smaller teams tend to disrupt science and technology with new ideas and opportunities, larger teams often concentrate on refining and developing existing concepts (Wuchty et al., 2007; Fortunato et al., 2018; Wu et al., 2019).
Effects of Discussion Turns on Novelty The number of discussion turns plays a crucial role in enhancing the novelty score (Mezirow, 2003; Li et al., 2023; Shinn et al., 2024; Lu et al., 2024). Our analysis (Fig. 3) indicates that an appropriate number of turns enables team members to explore topics thoroughly, iterate on ideas, and refine their abstracts. This iterative process is essential for deepening understanding and producing more sophisticated research outputs. While initial turns are beneficial for generating ideas and facilitating discussions, an excessive number of turns can lead to fatigue and reduced engagement. This may stifle creativity, as members may conform to dominant ideas rather than proposing innovative thoughts. Therefore, balancing the number of discussion turns is critical for maintaining high levels of novelty. Our findings suggest that peak novelty is achieved with a discussion turn count of 5.
Interaction Between Team Size and Discussion Turns Our findings further illustrate the interaction between team size and discussion turns. Larger teams with fewer turns can still produce relatively higher novelty scores, while smaller teams with excessive discussion turns may struggle to achieve the same level of innovation. This interplay emphasizes the importance of strategic planning in collaborative settings. Optimizing both team size and the number of discussion turns can significantly enhance the likelihood of generating novel outputs, indicating that the design of collaborative processes should take these factors into account.
4.3.2 How Team Freshness Affects Novelty
Figure 4: The balance of new and returning collaborators in the team has a notable impact on novelty, with 50% freshness yielding the highest historical dissimilarity and overall novelty, particularly in larger teams (size 8).
As shown in Fig. 4, team freshness, the fraction of team members who have not previously collaborated, has a notable effect on the novelty of generated outputs. Notably, team freshness shows its strongest effects at 50%, particularly for larger teams (size 8). At this level, historical dissimilarity reaches its peak, suggesting that a balanced mix of new and returning collaborators promotes divergence from past research, enhancing overall innovation. As team freshness increases, contemporary dissimilarity decreases, indicating that teams with fresh members tend to generate abstracts that align more closely with future research trends. Furthermore, both CI and ON achieve their highest values at 50% freshness. This suggests that a balanced team composition, where half the members are new, optimally combines novelty and future relevance, driving impactful research outcomes. Although our findings focus specifically on novelty in scientific abstracts, the broader principle aligns, to some extent, with prior work in this field (Guimera et al., 2005; Zeng et al., 2021).
4.3.3 How Team Research Diversity Affect Novelty
Figure 5: Effects of team diversity on novelty. The optimal diversity level appears to be 50%, which maximizes novelty and impact across team sizes.
Team research diversity, defined as the proportion of team members who specialize in unique research topics different from others in the team, plays a significant role in influencing various novelty metrics. As shown in Fig. 5, increasing diversity enhances HD, with both team sizes showing a peak at 50% diversity, indicating that moderately diverse teams produce work that is distinct from previous research. CD decreases most notably at 50% diversity, suggesting that teams with balanced diversity align better with future research trends while maintaining innovation. In terms of contemporary impact, larger teams (size 8) benefit more from higher diversity, seeing a significant increase, while smaller teams exhibit a more stable, moderate impact. Finally, Overall Novelty is highest at 50% diversity, particularly for larger teams, reflecting the value of having a balanced mix of diverse and non-diverse members for producing novel and impactful work. This conclusion mirrors findings from the Science of Science, where an unexpected combination of team members can increase research impact (Uzzi et al., 2013; Shi & Evans, 2023).
4.4 Ablation Study
Effects of Components Designed for Novelty In this section, we respectively test the effects of the invitation mechanism in team discussion, the role of the novelty assessment step, and the impact of self-review in abstract generation. All experiments are conducted with a 5-turn discussion. The results consistently show improvements in ON when these components are applied. For the invitation mechanism (Tab. 4.4), introducing new scientists into the discussion positively impacts the team’s performance across both 4-member and 8-member teams. This indicates that seeking external insights from relevant but non-team scientists fosters more diverse and novel ideas. The novelty assessment step (Tab. 4.4) also significantly boosts the scores. If novelty assessment is not considered, then the output of idea generation will not be an idea list but the idea from the last scientist. Novelty assessment ensures that the generated ideas are continuously evaluated for originality, helping teams avoid overlap with existing research. The improvement is most noticeable in larger teams, where more ideas are being generated and assessed. Finally, the self-review mechanism (Tab. 4.4) is crucial in further refining the abstracts. By allowing the team leader to re-evaluate the abstract for novelty after it is fully generated, low-quality abstracts are discarded, and the team engages in a new discussion to generate a better idea, as evidenced by the score improvements for both team sizes.
Table 2: Effects of invitation mechanism.
Size Invitation ON $\uparrow$
4 - 3.30
3.40
8 - 4.12
4.23
Table 3: Effects of novelty assessment.
Size Novelty Assessment ON $\uparrow$
4 - 3.19
3.40
8 - 3.98
4.23
Table 4: Effects of self-review in abstract generation.
Size Self-review ON $\uparrow$
4 - 3.26
3.40
8 - 3.99
4.23
Table 5: Comparison between fixed turns and adaptive turns in team discussions. The adaptive pattern shows both a lower inference cost and a higher ON.
Size Pattern Turns Inference Cost1 ON $\uparrow$
Topic Idea Check Abstract
4 Fixed 5 5 5 5 80 3.26
Adaptive 2.4 4.5 3.2 4.2 57.2 3.49
8 Fixed 5 5 5 5 160 3.99
Adaptive 2.9 4.8 4.0 3.8 124.0 4.37
  • 1
    The time taken by collaborator selection and team discussion’s invitation mechanism is not counted, and self-review is not employed for a better comparison.
Effects of Discussion Pattern In the previous experiments, we fixed the number of discussion turns in each step to ensure fair comparisons. However, in real-world research environments, teams of scientists do not spend the same amount of time on each stage of the research process. To explore this, we compare fixed discussion turns with adaptive turn numbers. In the adaptive pattern, the team leader decides whether the team needs additional turns based on the current progress and the goals of each stage. The results of both patterns, along with their corresponding inference cost, are shown in Tab. 5. The comparison reveals that the adaptive pattern achieves a higher ON while reducing the inference cost. This efficiency can be attributed to the more flexible approach, allowing teams to adjust their discussions dynamically rather than adhering to a rigid structure. Furthermore, examining the number of turns at each stage in both 4-person and 8-person teams under the adaptive pattern offers additional insights. Larger teams require more discussion turns and face greater challenges in reaching consensus (Janis, 1972; Pitters & Oberlechner, 2014). This highlights the adaptive pattern’s advantage in accommodating the complexities of larger teams while maintaining a higher level of novelty in the final research outputs.
5 Conclusion
We build the virtual scientist (VirSci), a pioneering LLM-based multi-agent system designed to replicate the collaborative dynamics of scientific discovery. By addressing the limitations of traditional single-agent systems, our model effectively simulates the initial phase of autonomous science discovery—idea generation. Through a structured five-step process, VirSci showcases how specialized agents can collaborate, offering diverse expertise and insights that mirror real-world scientific teamwork. The experiments reveal that VirSci significantly outperforms single-agent approaches and highlights emergent social behaviors among the scientist agents, suggesting promising avenues for further exploration of collaborative mechanisms in scientific research.
Appendix A Data and Code Availability
A.1 Code
The code for this project is available at https://github.com/open-sciencelab/Social_Science.
A.2 Data
Appendix B Limitations and Future Work
While our multi-agent system demonstrates notable improvements over the single-agent approach in generating novel and impactful scientific ideas, it has several limitations. First, we have only validated our system on a single computer science dataset, which restricts the diversity of research ideas and limits its capacity for simulating interdisciplinary collaborations. This focus on a single domain also reduces the generalizability of our results to other scientific fields. Additionally, while our system effectively models collaboration, the simulated interactions may oversimplify the complexities of real-world teamwork, where multiple teams can collaborate dependently or independently on related research, and agents often participate in different teams simultaneously.
To address these limitations, several future directions can be pursued. Expanding the system to incorporate datasets from various scientific disciplines is a crucial next step. This would increase the diversity of generated ideas and enable simulations of interdisciplinary collaborations, providing a more realistic and holistic representation of real-world research environments. Another important direction is to enhance the simulation of teamwork by allowing multiple teams to work concurrently and enabling agents to contribute to multiple teams or projects simultaneously. Such improvements would better reflect the collaborative dynamics of modern scientific research and provide a more powerful tool for the Science of Science community. This would allow researchers to probe deeper into the underlying processes of scientific collaboration, engaging with the dynamic and interactive nature of teamwork to gain insights into how collaboration fosters innovation.
Appendix C Ethics Statement
This research uses publicly available data from the AMiner dataset, ensuring compliance with data privacy policies. Author names are masked to prevent data leakage during simulations. Our system is intended to augment, not replace, human researchers, emphasizing the need for human oversight to ensure the quality and integrity of generated outputs. To promote transparency, we commit to sharing all relevant codes for reproducibility within the research community.
Appendix D Effect of the Potential Data Leakage
We acknowledge that the use of papers published between 2011 and 2014 may raise concerns about data leakage, given that the LLMs employed in our experiments are trained on data within this time period. However, this potential issue does not pose a significant threat to the validity of our experiments for the following reasons. First, both the comparisons between our multi-agent system and AI Scientist, as well as the comparisons between different team settings, utilize the same LLMs. Since all models encounter the same exposure to training data from this period, any potential data leakage would affect all experiments equally. Thus, the relative performance differences we observe are not skewed by uneven data leakage. This ensures that the evaluation process remains fair and that the corresponding conclusions drawn are valid. Moreover, our goal is not to demonstrate an absolute measure of novelty but rather to explore how different collaboration strategies and team settings influence the novelty of generated research outputs. As all team settings face the same potential exposure to historical data, the novelty metrics still provide an accurate comparison of the agents’ ability to generate distinct and original ideas under varying conditions. In summary, while data leakage is a valid concern, it affects all models and settings uniformly in our experiments. Therefore, it does not undermine the relative comparisons we make or the conclusions we draw regarding collaboration strategies and team performance.
Appendix E More Details of Methods
E.1 Self-review
A self-review mechanism is considered after $R_{abstract}$ is finalized to pre-check its novelty. In this self-review, the optimized abstract $R_{abstract}$ is provided to the team leader to assess novelty by comparing it to similar papers in $B_{past}$. The prompt is:
$Q_{review}=\langle Q_{check},R_{abstract},B_{past}(R_{abstract})\rangle$ (7)
If this is the first time undergoing the self-review and the team leader determines that the similarity to existing papers is too high, the abstract will undergo further revision. The evaluation $R_{review}$ will then be added to Eq. (6) for the next revision round:
$Q_{1,0}=\langle Q_{abstract},Q_{judgement},R_{review},B_{past}(R_{abstract}),R_{abstract}\rangle$ (8)
If the abstract undergoes a second self-review and still does not meet the novelty requirement, it will be discarded, and the team will generate a new idea. Once the self-review yields satisfactory results, the final abstract will be produced, and the system will terminate. However, this self-review mechanism introduces uncertainty in total inference cost, making it difficult to ensure fair experimental comparisons. We discuss the effectiveness of this module only in the ablation study (see Sec. 4.4).
Appendix F More Experiments
To evaluate the validity of our proposed overall novelty metric, We extract a total of 100 abstracts generated under different experimental conditions and have them evaluated by (1) our overall novelty metric, (2) an LLM-based reviewer (where we use the GPT-4o API with version “gpt-4o-2024-08-06”), and (3) human researchers in the computer science domain. For the LLM-based reviewer and human researchers, we use the idea review form in (Si et al., 2024) for scoring. The evaluation results are presented in Fig. 6 and Fig. 7, where the axes of Fig. 6 represent the score of the same abstract evaluated under different metrics for (1) and (2), and the axes of Fig 7 represent the score of the same abstract evaluated under different metrics for (1) and (3). The Pearson correlation coefficients between our proposed overall novelty metric and LLM-based reviewers, as well as between our metric and human researchers, demonstrate the positive correlation of our metric with currently used novelty measurement methods (Lu et al., 2024; Si et al., 2024), which, to some extent, supports the validity of our metric.
Figure 6: The evaluation results of the same abstract under two different review metrics: our proposed overall novelty metric and LLM-based reviewer. The Pearson correlation coefficient equals 0.59, denoting the positive correlation of our metric with the LLM-based reviewer.
Figure 7: The evaluation results of the same abstract under two different review metrics: our proposed overall novelty metric and human researcher. The Pearson correlation coefficient equals 0.51, denoting the positive correlation of our metric with the human researcher.
Appendix G Prompts
G.1 Scientist Definition
We use the personal information of the scientist to define the agent, where the corresponding system prompt is illustrated in Fig. 8.
Figure 8: The system prompt of each scientist agent is the personal information, including the name, role, affiliation, research interests, citation situation, and collaboration history.
G.2 Collaboration Selection
The prompt for collaboration selection is illustrated in Fig. 9.
Figure 9: The prompt for the collaboration selection.
G.3 Topic Discussion
G.3.1 Discussion
The prompt for the topic discussion is illustrated in Fig. 10.
Figure 10: The prompt for the topic discussion.
G.3.2 Summarization
The prompt for the final topic selection after several turns of topic discussion is illustrated in Fig. 11.
Figure 11: The prompt for the final topic selection after topic discussion.
G.4 Idea Generation
The prompt for the idea generation is illustrated in Fig. 12.
Figure 12: The prompt for the idea generation.
G.5 Novelty Assessment
The prompt for the novelty assessment is illustrated in Fig. 13.
Figure 13: The prompt for the novelty assessment.
G.6 Abstract Generation
G.6.1 Discussion
The prompt for the beginning case of the abstract generation is illustrated in Fig. 14.
Figure 14: The prompt for the beginning case of the abstract generation.
The prompt for the normal case of the abstract generation is illustrated in Fig. 15.
Figure 15: The prompt for the normal case of the abstract generation.
G.6.2 Self-review
The prompt for the self-review after generating the final abstract is illustrated in Fig. 16.
Figure 16: The prompt for the self-review after generating the final abstract.
G.7 LLM Review
The prompt for the LLM-based review is based on NeurIPS2024 reviewer guidelines, which is the same metric as AI Scientist to ensure a fair comparison between our method and AI Scientist. The content is illustrated in Fig. 17.
Figure 17: LLM review. To ensure a fair comparison, we use the same metric as AI Scientist, which is based on NeurIPS2024 reviewer guidelines. We only keep several critical metrics in this guideline since now we only need to evaluate the abstract.
Appendix H Example Scenarios
H.1 Collaboration Selection
The example scenario of the collaborator selection is illustrated in Fig. 18. Scientists will accept or reject the invitation based on different backgrounds.
Figure 18: The example scenario of the collaborator selection. Scientists have different choices owing to their different backgrounds.
H.2 Topic Discussion
H.2.1 Topic Discussion Normal Case
The example scenario of the normal case in the topic discussion is illustrated in Fig. 19. Scientists will provide topic discussion responses, which ensure a coherent discussion of the research topic.
Figure 19: The example scenario of the normal case in the topic discussion. Scientists will provide topic discussion responses (highlighted in yellow), which ensure a coherent discussion of the research topic.
H.2.2 Invitation Mechanism
The example scenario of the invitation mechanism in the topic discussion is illustrated in Fig. 20, which ensures a comprehensive topic discussion.
Figure 20: The example scenario of the invitation mechanism in the topic discussion. We highlight the content of the collaboration invitation mechanism in blue.
H.3 Idea Generation
The example scenario of the beginning case of the idea generation is illustrated in Fig. 21.
Figure 21: The example scenario of the beginning case of the idea generation.
The example scenario of the normal case in the idea generation is illustrated in Fig. 22.
Figure 22: The example scenario of the normal case in the idea generation.
H.4 Novelty Assessment
The example scenario of the user prompt provided for scientist agents in the novelty assessment is illustrated in Fig. 23. The prompt includes three candidate ideas and related papers.
Figure 23: The example scenario of the user prompt provided for scientist agents in the novelty assessment. There are three ideas and related papers.
The example scenario of the agent responses in the novelty assessment is illustrated in Fig. 24. Note that Fig. 24 corresponds to Fig. 23.
Figure 24: The example scenario of the agent responses in the novelty assessment. By max-voting, idea 2 is selected as the final idea.
H.5 Abstract Generation
H.5.1 Abstract Generation Normal Case
The example scenario of the beginning case in the abstract generation is illustrated in Fig. 25.
Figure 25: The example scenario of the beginning case in the abstract generation.
The example scenario of the normal case in the abstract generation is illustrated in Fig. 26.
Figure 26: The example scenario of the normal case in the abstract generation.
H.5.2 Self-review
The example scenario of the self-review results is illustrated in Fig. 27.
Figure 27: The example scenario of the self-review results.